Strategic placement of lobster traps is one of the keys for a successful lobster fisherman. A study was conducted of the average distance separating traps long dash—called trap spacinglong dash—deployed by lobster fishermen. The​ trap-spacing measurements​ (in meters) for a sample of seven teams from fishing cooperative​ A, and a sample of eight teams from fishing cooperative B are repeated in the accompanying table. For this​ problem, we are interested in comparing the mean​ trap-spacing measurements of the two fishing cooperatives.​ (Since the sample from fishing cooperative A is listed​ first, treat it as the first​ sample.)

Find a 90% confidence interval for the target parameter.

The confidence interval is ( _ , _ )

(Round to one decimal as needed.)

Please try and show work. THANK YOU!

Table of Trap Spacing Measurements: The left side of the table is A cooperative (total # = 7) and the right side is B cooperative (total # = 8).

80 152
90 103
85 95
73 105
80 69
73 102
83 66
116
Answer #1

Population parameter is the difference between mean trap spacing measurements of two fishing cooperatives that is \mu 1-\mu 1

90% confidence interval for \mu 1-\mu 1 is

(21 – 22) Ete** 1/n1+1/n2)

cl = 80.57, r2 = 101, s12 = 38.2857, s2 = 733.1429, n1 = 7, n2 = 8

s = V(nl - 1)s12 + (n 2 - 1)s22)/(n1 + n2-2)

=20.3086

degrees of freedom = n1+n2-2 = 7+8-2=13

For 90% confidence(a = 0.10) with df =13 , two tailed critical value of t is

tc = 1.77 (from t table)

Thus,90% confidence interval for \mu 1-\mu 1 is

(80.57 – 101) + 1.77 * 20.3086 * (1/7 +1/8)

=(-20.435+18.60

=(- 39.0 , -1.8)

Calculation :

x1 (x1-\bar{x1})^2 x2 (x2-\bar{x2})^2
80 0.3249 152 2601
90 88.9249 103 4
85 19.6249 95 36
73 57.3049 105 16
80 0.3249 69 1024
73 57.3049 102 1
83 5.9049 66 1225
sum/ss 564 229.7143 116 225
\bar{x1}=sum/7 80.57143 sum/ss 808 5132
s1^2=ss/6 38.28571667 \bar{x2}=sum/8 101
s2^2=ss/7 733.1428571